[figure: 9.7 cm]

SEARCH FOR CONDITIONS OF CO-EXISTENCE OF TAMM AND SHOCKLEY STATES IN SUPERLATTICE WITH PERTURBED SURFACE

Jarosław Kłos and Henryk Puszkarski Surface Physics Division, Faculty of Physics, Adam Mickiewicz University, ul. Umultowska 85, 61-614 Poznań, Poland

This study is based on the surface states classification proposed by Zak[1], and the analysis is performed using the transfer matrix method as formulated by James[2]. Conditions for existence of electron surface states are studied in a semi-infinite superlattice with alloy concentration changing either abruptly (square potential), or gradually (sinusoidal potential). A small surface perturbation is introduced into this model by modifying the shape of a potential well near the surface. The effect of this surface perturbation on Shockley states is studied, in order to find the threshold perturbation which destroys the existing surface state, shifting it from the gap to a band. For Tamm states, we search for a minimum perturbation inducing surface states to appear. The results obtained do show co-existence of Tamm and Shockley states within the specific range of the variable model parameters.

[1] J. Zak, Physical Review B32, 2218 (1985).

[2] H.M. James, Physical Review 76, 1602 (1949).

[figure: 13.4 cm]

Subject category :

7. Computations of Electronic Band Structures

Presentation mode : poster

Corresponding author : Jarosław Kłos

Address for correspondence : Surface Physics Division, Faculty of Physics, Adam Mickiewicz University, ul. Umultowska 85, 61-614 Poznań, Poland

Email address : klos@amu.edu.pl